# MonoMAE: Enhancing Monocular 3D Detection through Depth-Aware Masked Autoencoders

### v1

@@ Abstract
Monocular 3D object detection aims for precise 3D localization and identification of objects from a single-view image. Despite its recent progress, it often struggles while handling pervasive object occlusions that tend to complicate and degrade the prediction of object dimensions, depths, and orientations. We design MonoMAE, a monocular 3D detector inspired by Masked Autoencoders that addresses the object occlusion issue by masking and reconstructing objects in the feature space. MonoMAE consists of two novel designs. The first is depth-aware masking that selectively masks certain parts of non-occluded object queries in the feature space for simulating occluded object queries for network training. It masks non-occluded object queries by balancing the masked and preserved query portions adaptively according to the depth information. The second is lightweight query completion that works with the depth-aware masking to learn to reconstruct and complete the masked object queries. With the proposed object occlusion and completion, MonoMAE learns enriched 3D representations that achieve superior monocular 3D detection performance qualitatively and quantitatively for both occluded and non-occluded objects. Additionally, MonoMAE learns generalizable representations that can work well in new domains.
Keywords:
Monocular 3D Object Detection Masked Autoencoders Occlusion
@@ 1 Introduction
Figure 1: Object occlusion is pervasive and affects monocular 3D detection: Object occlusion is pervasive, e.g., 62% (17725) cars in the KITTI 3D dataset suffer from various occlusions as illustrated in (a). Prevalent monocular 3D detection techniques such as GUPNet [31] and MonoDETR [57] are clearly affected by object occlusions in both 3D space (3D) and the bird’s eye view (BEV) space as in (b). Our MonoMAE simulates and learns object occlusions by feature masking and completing which improves detection consistently for both occluded and non-occluded objects.
@@
Among various new challenges in monocular 3D detection, object occlusion, which exists widely in natural images as illustrated in Figure 1 (a), becomes a critical issue while predicting 3D location in terms of object depths, object dimensions, and object orientations. Most existing monocular 3D detectors such as MonoDETR[57] and GUPNet [31] neglect the object occlusion issue which demonstrates clear performance degradation as illustrated in Figure 1 (b). A simple idea is to learn to reconstruct the occluded object regions whereby occluded objects can be handled similarly as non-occluded objects. On the other hand, reconstructing occluded object regions in the image space is complicated due to the super-rich variation of object occlusions in scene images.
@@
Inspired by the Masked Autoencoders (MAE) [15] that randomly occludes image patches and reconstructs them in representation learning, we treat object occlusions as natural masking and train networks to complete occluded object regions to learn occlusion-tolerant representations. To this end, we design MonoMAE, a novel monocular 3D detection framework that adopts the idea of MAE by first masking certain object regions in the feature space (for simulating object occlusions) and then reconstructing the masked object features (for learning occlusion-tolerant representations). MonoMAE consists of a depth-aware masking module and a lightweight completion network. The depth-aware masking simulates object occlusions by masking the features of non-occluded objects adaptively according to the object depth information. It generates pairs of non-occluded and masked (i.e., occluded) object representations that can be directly applied to train the lightweight completion network, aiming for completing the occluded objects and learning occlusion-tolerant representations. Note that MonoMAE introduces little computational overhead in inference time as it requires no object masking in the inference stage and the completion network is lightweight. Moreover, completion in the feature space can be effectively achieved using a lightweight network, surpassing the complex encoder-decoder structure required for image space completion.
@@
The contributions of this work can be summarized in three major aspects. First, we design MonoMAE, a MAE-inspired monocular 3D detection framework that tackles object occlusions effectively by masking and reconstructing object regions at the feature level. Second, we design adaptive image masking and a lightweight completion network that mask non-occluded objects adaptively according to the object depth (for simulating object occlusions) and reconstruct the masked object (for learning occlusion-tolerant representations), respectively. Third, extensive experiments over KITTI 3D and nuScenes show that MonoMAE outperforms the state-of-the-art consistently and it can generalize to new domains as well.
@@ 2.1 Monocular 3D Object Detection
Monocular 3D detection aims for identification and 3D localization of objects from a single-view image. Most existing work can be broadly classified into two categories. The first employs convolutional neural networks and most methods adopt center-guided pipelines following conventional 2D detectors [12]. The standard approach learns monocular 3D detectors from single-view images only [17, 1, 8, 61, 57]. To acquire more depth information, several studies explore to leverage extra training data, such as LiDAR point clouds [33, 48, 6, 41], depth maps [11, 40, 39, 21], and 3D CAD models [7, 29, 36]. Beyond that, several studies exploit the geometry relation between 2D and 3D spaces in different ways. For example, M3D-RPN [1] applies the powerful 2D detector FPN [42] for 3D detection. MonoDLE [34] aligns the centers of 2D and 3D boxes for better 3D localization. GUPNet [31] leverages uncertainty modeling to estimate the height of 3D boxes from the 2D boxes.
The second introduces powerful visual transformers for more accurate monocular 3D detection  [19, 57, 62, 50, 49]. For example, MonoDTR [19] integrates context- and depth-aware features and injects depth positional hints into transformers. MonoDETR [57] modifies the transformer to be depth-aware and guides the detection process by contextual depth cues. However, most existing studies neglect object occlusions that exist widely in natural images and often degrade the performance of monocular 3D object detection clearly. We adopt the transformer architecture to learn occlusion-tolerant representations that can handle object occlusion effectively without requiring any extra data.
@@ 2.2 Occlusions in 3D Object Detection
Object occlusion is pervasive in scene images and it has been investigated in several 3D detection studies [52, 25, 18, 8, 55, 45, 10]. One typical approach learns to estimate the complete localization of occluded objects. For example, Mono-3DT [18] estimates complete 3D bounding boxes by re-identifying occluded vehicles from a sequence of 2D images. BtcDet [52] leverages object shape priors and learns to estimate the complete shapes of partially occluded objects. Several studies consider the degree of occlusions in training. For example, MonoPair [8] exploits the relation of paired samples and encodes spatial constraints of occluded objects from their neighbors. HMF [25] introduces an anti-occlusion loss to focus on occluded samples. Different from previous methods, our MonoMAE learns enriched and occlusion-tolerant representations by masking and completing object parts in the feature space.
@@ 2.3 Masked Autoencoders in 3D Tasks
Masked Autoencoders (MAE) [15] learn visual representations by masking image patches and reconstructing them, and it has been explored in several point cloud pre-training studies. For outdoor point cloud pre-training, Occupancy-MAE [35] exploits range-aware random masking that employs three masking levels to deal with the sparse voxel occupancy structures of LiDAR point clouds. GD-MAE [54] introduces a Generative Decoder to merge the surrounding context to restore the masked tokens hierarchically. For indoor point cloud pre-training, Point-MAE [37] adopts MAE to directly reconstruct the 3D coordinates of masked tokens. I2P-MAE [58] introduces 2D pre-trained models which enhances 3D pre-training with diverse 2D semantics. PiMAE [5] learns cross-modal representations with MAE by interactively handling point clouds and RGB images. Different from the above studies, our MonoMAE handles monocular 3D detection from single-view images and it focuses on object occlusions by learning to complete occluded object regions at the feature level.
@@ 3.1 Problem Definition
Monocular 3D detection takes a single RGB image as input, aiming to classify objects and predict their 3D bounding boxes. The prediction of each object is composed of the object category $C$, a 2D bounding box $B_{2D}$, and a 3D bounding box $B_{3D}$ where $B_{3D}$ can be further decomposed to the object 3D location $(x_{3D},y_{3D},z_{3D})$, the object dimensions in object height, width and length $(h_{3D},w_{3D},l_{3D})$, as well as the object orientation $\theta$.
Figure 2: The training framework: Given a single-view image, the 3D Backbone extracts 3D object query features which are grouped into non-occluded query features and occluded query features by the Non-Occluded Query Grouping. The Depth-Aware Masking then masks the non-occluded query features to simulate object occlusions according to the object depth, and the Completion Network further learns to reconstruct the masked queries. Finally, the completed and the occluded query features are concatenated to train the 3D Detection Head for 3D predictions.
Figure 3: The inference framework. Given a test image, the 3D Backbone extracts 3D object queries and feeds the queries to the Non-Occluded Query Grouping for classification. The identified occluded queries are then fed to the Completion Network to reconstruct the completed queries. Finally, the completed queries and the identified non-occluded queries are concatenated and fed to the 3D Detection Head for 3D detection predictions.
@@ 3.2 Overall Framework
Figure 2 shows the framework of the proposed MonoMAE. Given an input image $I$, the 3D Backbone first generates a sequence of 3D object queries $Q=[q_{1},q_{2},\cdots,q_{K}]$ ($K$ denotes query number) and the Non-Occluded Query Grouping then classifies the queries into two groups including non-occluded queries $Q^{NO}=[q_{1}^{NO},q_{2}^{NO},\cdots,q_{U}^{NO}]$ and occluded queries $Q^{O}=[q_{1}^{O},q_{2}^{O},\cdots,q_{V}^{O}]$ ($U$ and $V$ are the number of non-occluded and occluded queries). The Non-Occluded Query Masking then masks $Q^{NO}$ to produce masked queries according to their depth $D=[d_{1},d_{2},\cdots,d_{U}]$, leading to the masked queries $Q^{M}=[q_{1}^{M},q_{2}^{M},\cdots,q_{U}^{M}]$. The Query Completion further reconstructs $Q^{M}$ to produce the completed queries $Q^{C}=[q_{1}^{C},q_{2}^{C},\cdots,q_{U}^{C}]$. Finally, the occluded queries $Q^{O}$ and the completed queries $Q^{C}$ are concatenated and fed to the Monocular 3D Detection for 3D detection predictions. Note the inference does not involve the Non-Occluded Query Masking, and it just concatenates the completion of occluded queries $Q^{O}$ (i.e., $Q^{C}$) with the non-occluded queries $Q^{NO}$ and feeds the concatenated queries to the 3D Detection Head for 3D predictions as illustrated in Figure 3.
@@ 3.3 Non-Occluded Query Masking
Queries predicted by the 3D Backbone are either occluded or non-occluded, depending on whether the corresponding objects are occluded in the input image. In MonoMAE, we mask the non-occluded queries in the feature space to simulate occlusions, aiming to generate pairs of non-occluded and masked (i.e., occluded) queries for learning occlusion-tolerant object representations. Specifically, we design Non-Occluded Query Grouping to identify non-occluded queries and feed them into the Depth-Aware Masking to synthesize occlusions.
Specifically, we design a Non-Occluded Query Grouping module to identify non-occluded queries and then feed them into a designed Depth-Aware Masking module to synthesize occlusions, with more detail to be elaborated in the following subsections.
Non-Occluded Query Grouping. The Non-Occluded Query Grouping classifies the queries based on whether their corresponding objects are occluded or non-occluded. With no information about whether the input queries are occluded, we design an occlusion classification network $\mathrm{\Phi}_{\text{O}}$ to predict the occlusion condition $O^{p}=[o_{1}^{p},o_{2}^{p},\cdots,o_{K}^{p}]$ of queries $Q=[q_{1},q_{2},\cdots,q_{K}]$, where for the $i$-th query $o_{i}^{p}=\mathrm{\Phi}_{\text{O}}(q_{i})$. The Non-Occluded Query Grouping can be formulated by:
@@
where $CE$ is the $Cross$ $Entropy$ loss. We adopted the bipartite matching [4] to match the predicted queries and objects in the image, where only matched queries have ground truth $O^{gt}$ of KITTI 3D [13] about whether they are occluded or not. Please refer to the supplementary materials for details.
Figure 4: Illustration of the Depth-Aware Masking. (a) Objects farther away are usually smaller with less visual information. (b) The Depth-Aware Masking determines the mask ratio of an object according to its depth - the closer the object is, the larger the mask ratio is applied, thereby compensating the information deficiency for objects that are distant from the camera.
Depth-Aware Masking. We design depth-aware masking to adaptively mask non-occluded query features to simulate occlusions at the feature level, aiming to create non-occluded and occluded (i.e., masked) pairs for learning occlusion-tolerant representations. As illustrated in Figure 4, the depth-aware masking determines the mask ratio according to the object depth - the closer the object, the larger the mask ratio, thereby compensating the information deficiency of distant objects. In addition, we simulate occlusions by masking at the feature level, as masking and reconstructing at the image level is complicated and computationally intensive.
@@
where $q_{i}^{M}$ is the $i$-th masked query, $q_{i}^{NO}$ is the $i$-th non-occluded query, and $m_{i}$ is the generated mask.
@@ 3.4 Query Completion
The query completion learns to reconstruct the adaptively masked queries, aiming to produce completed queries whereby the network learns occlusion-tolerant representations that are helpful in detecting occluded objects. We design a completion network $\mathrm{\Phi}_{\text{C}}$ to reconstruct the masked queries. The Completion Network has an hourglass structure consisting of three conv-bn-relu blocks and one conv-bn block for 3D query completion. Please refer to the appendix for details. The completed query $q_{i}^{C}$ is obtained by:
@@
3.5 Loss Functions
The overall objective consists of three losses including $L_{occ}$, $L_{com}$, and $L_{base}$ where $L_{occ}$ and $L_{com}$ are defined in Equation 2 and Equation 6, and $L_{base}$ denote losses for supervising the 3D box predictions.
$L_{base}$ includes losses for supervising the 3D box predictions including each object’s 3D locations, height, width, length and orientation. We set the weight for each loss item to 1.0, and the overall loss function is formulated as follows:
$L=L_{occ}+L_{com}+L_{base}.$ (7)
4 Experiments
4.1 Experimental Settings
$\bullet$ NuScenes [3] comprises 1,000 video scenes, including RGB images captured by 6 surround-view cameras. The dataset is split into a training set (700 scenes), a validation set (150 scenes), and a test set (150 scenes).
Following [1, 43, 31, 22, 20], the performance on the validation set is reported. We follow prior studies [44, 43, 47, 57] that evaluate only on the most representative Car category of KITTI 3D and nuScenes datasets.
[…]
Implementation Details. We conduct experiments on one NVIDIA V100 GPU and train the framework for 200 epochs with a batch size of 16 and a learning rate of $2\times 10^{-4}$. We use the AdamW [30] optimizer with weight decay $10^{-4}$. We employ ResNet-50 [16] as the Transformer-based backbone and adopt the 3D detection head from [57] for our framework.
@@ 4.1 Experimental Settings
PatchNet [32] ECCV 20 Depth 15.68 11.12 10.17 22.97 16.86 14.97
D4LCN [11] CVPR 20 16.65 11.72 9.51 22.51 16.02 12.55
@@
M3D-RPN [1] ICCV 19 None 14.76 9.71 7.42 21.02 13.67 10.23
SMOKE [28] CVPRW 20 14.03 9.76 7.84 20.83 14.49 12.75
MonoPair [8] CVPR 20 13.04 9.99 8.65 19.28 14.83 12.89
MonoDLE [34] CVPR 21 17.23 12.26 10.29 24.79 18.89 16.00
MonoFlex [59] CVPR 21 19.94 13.89 12.07 28.23 19.75 16.89

### v2

@@ Abstract
Monocular 3D object detection aims for precise 3D localization and identification of objects from a single-view image. Despite its recent progress, it often struggles while handling pervasive object occlusions that tend to complicate and degrade the prediction of object dimensions, depths, and orientations. We design MonoMAE, a monocular 3D detector inspired by Masked Autoencoders that addresses the object occlusion issue by masking and reconstructing objects in the feature space. MonoMAE consists of two novel designs. The first is depth-aware masking that selectively masks certain parts of non-occluded object queries in the feature space for simulating occluded object queries for network training. It masks non-occluded object queries by balancing the masked and preserved query portions adaptively according to the depth information. The second is lightweight query completion that works with the depth-aware masking to learn to reconstruct and complete the masked object queries. With the proposed feature-space occlusion and completion, MonoMAE learns enriched 3D representations that achieve superior monocular 3D detection performance qualitatively and quantitatively for both occluded and non-occluded objects. Additionally, MonoMAE learns generalizable representations that can work well in new domains.
@@ 1 Introduction
Figure 1: Object occlusion is pervasive and affects monocular 3D detection: Object occlusion is pervasive, e.g., 62% (17725) cars in the KITTI 3D dataset suffer from various occlusions as illustrated in (a). Prevalent monocular 3D detection techniques such as GUPNet [37] and MonoDETR [66] are clearly affected by object occlusions in both 3D space (3D) and the bird’s eye view (BEV) space as in (b). The proposed MonoMAE simulates and learns object occlusions by feature masking and completing which improves detection consistently for both occluded and non-occluded objects.
@@
Among various new challenges in monocular 3D detection, object occlusion, which exists widely in natural images as illustrated in Figure 1 (a), becomes a critical issue while predicting 3D locations in terms of object depths, object dimensions, and object orientations. Most existing monocular 3D detectors such as MonoDETR[66] and GUPNet [37] neglect the object occlusion issue which demonstrates clear performance degradation as illustrated in Figure 1 (b). A simple idea is to learn to reconstruct the occluded object regions whereby occluded objects can be handled similarly as non-occluded objects. On the other hand, reconstructing occluded object regions in the image space is complicated due to the super-rich variation of object occlusions in scene images.
@@
Inspired by the Masked Autoencoders (MAE) [15] that randomly occludes image patches and reconstructs them in representation learning, we treat object occlusions as natural masking and train networks to complete occluded object regions to learn occlusion-tolerant representations. To this end, we design MonoMAE, a novel monocular 3D detection framework that adopts the idea of MAE by first masking certain object regions in the feature space (for simulating object occlusions) and then reconstructing the masked object features (for learning occlusion-tolerant representations). MonoMAE consists of a depth-aware masking module and a lightweight completion network. The depth-aware masking simulates object occlusions by masking the features of non-occluded objects adaptively according to the object depth information. It generates pairs of non-occluded and masked (i.e., occluded) object representations that can be directly applied to train the lightweight completion network, aiming for completing the occluded objects and learning occlusion-tolerant representations. Note that MonoMAE introduces little computational overhead in inference time as it requires no object masking in the inference stage.
@@
The contributions of this work can be summarized in three major aspects. First, we design MonoMAE, a MAE-inspired monocular 3D detection framework that tackles object occlusions effectively by masking and reconstructing object regions in the feature space. To the best of our knowledge, this is the first work that explores masking-reconstructing for the task of monocular 3D object detection. Second, we design adaptive image masking and a lightweight completion network that mask non-occluded objects adaptively according to the object depth (for simulating object occlusions) and reconstruct the masked object regions (for learning occlusion-tolerant representations), respectively. Third, extensive experiments over KITTI 3D and nuScenes show that MonoMAE outperforms the state-of-the-art consistently and it can generalize to new domains as well.
@@ 2.1 Monocular 3D Object Detection
Monocular 3D detection aims for the identification and 3D localization of objects from a single-view image. Most existing work can be broadly classified into two categories. The first employs convolutional neural networks, where most methods follow conventional 2D detectors [12, 23]. The standard approach learns monocular 3D detectors from single-view images only [17, 1, 8, 70, 66, 34]. In addition, several studies explore to leverage extra training data, such as LiDAR point clouds [39, 55, 6, 48, 46, 45], depth maps [11, 47, 45, 22, 38], and 3D CAD models [7, 35, 42] to acquire more depth information. Beyond that, several studies exploit the geometry relation between 2D and 3D spaces in different ways. For example, M3D-RPN [1] applies the powerful 2D detector FPN [49] for 3D detection. MonoDLE [40] aligns the centers of 2D and 3D boxes for better 3D localization. GUPNet [37] leverages uncertainty modeling to estimate the height of 3D boxes from the 2D boxes.
The second introduces powerful visual transformers [72, 21, 4, 65] for more accurate monocular 3D detection  [19, 66, 71, 57, 56]. For example, MonoDTR [19] integrates context- and depth-aware features and injects depth positional hints into transformers. MonoDETR [66] modifies the transformer to be depth-aware and guides the detection process by contextual depth cues. However, most existing methods neglect object occlusions that exist widely in natural images and often degrade the performance of monocular 3D object detection clearly. We adopt the transformer architecture to learn occlusion-tolerant representations that can handle object occlusion effectively without requiring any extra training data or annotations.
@@ 2.2 Occlusions in 3D Object Detection
Object occlusion is pervasive in scene images and it has been investigated in several 2D and 3D vision tasks [63, 52, 10, 28, 26, 29, 30]. One typical approach learns to estimate the complete localization of occluded objects. For example, Mono-3DT [18] estimates complete 3D bounding boxes by re-identifying occluded vehicles from a sequence of 2D images. BtcDet [59] leverages object shape priors to learns to estimate the complete shapes of partially occluded objects. Several studies consider the degree of occlusions in training. For example, MonoPair [8] exploits the relation of paired samples and encodes spatial constraints of occluded objects from their neighbors. HMF [31] introduces an anti-occlusion loss to focus on occluded samples. Different from existing methods, the proposed MonoMAE learns enriched and occlusion-tolerant representations by masking and completing object parts in the feature space.
@@ 2.3 Masked Autoencoders in 3D Tasks
Masked Autoencoders (MAE) [15] learn visual representations by masking image patches and reconstructing them, and it has been explored in several point cloud pre-training studies. For outdoor point cloud pre-training, Occupancy-MAE [41] exploits range-aware random masking that employs three masking levels to deal with the sparse voxel occupancy structures of LiDAR point clouds. GD-MAE [62] introduces a Generative Decoder to merge the surrounding context to restore the masked tokens hierarchically. For indoor point cloud pre-training, Point-MAE [43] adopts MAE to directly reconstruct the 3D coordinates of masked tokens. I2P-MAE [67] introduces 2D pre-trained models, and it enhances 3D pre-training with diverse 2D semantics. PiMAE [5] learns cross-modal representations with MAE by interactively handling point clouds and RGB images. Different from existing studies, the proposed MonoMAE handles monocular 3D detection from single-view images and it focuses on object occlusions by learning to complete occluded object regions in the feature space.
@@ 3.1 Problem Definition
Monocular 3D detection takes a single RGB image as input, aiming to classify objects and predict their 3D bounding boxes. The prediction of each object is composed of the object category $C$, a 2D bounding box $B_{2D}$, and a 3D bounding box $B_{3D}$. The 3D bounding box $B_{3D}$ can be decomposed to the object 3D location $(x_{3D},y_{3D},z_{3D})$, the object dimensions in object height, width and length $(h_{3D},w_{3D},l_{3D})$, as well as the object orientation $\theta$.
Figure 2: The framework of MonoMAE training: Given a single-view image, the 3D Backbone extracts 3D object query features which are grouped into non-occluded query features and occluded query features by the Non-Occluded Query Grouping. The Depth-Aware Masking then masks the non-occluded query features to simulate object occlusions adaptively based on the object depth, and the Completion Network then learns to reconstruct the masked queries. Finally, the completed and the occluded query features are concatenated to train the 3D Detection Head for 3D predictions.
@@ 3.2 Overall Framework
Figure 2 shows the framework of the proposed MonoMAE. Given an input image $I$, the 3D Backbone first generates a sequence of 3D object queries $Q=[q_{1},q_{2},\cdots,q_{K}]$ ($K$ denotes query number), and the Non-Occluded Query Grouping then classifies the queries into two groups including non-occluded queries $Q^{NO}=[q_{1}^{NO},q_{2}^{NO},\cdots,q_{U}^{NO}]$ and occluded queries $Q^{O}=[q_{1}^{O},q_{2}^{O},\cdots,q_{V}^{O}]$ ($U$ and $V$ are the number of non-occluded and occluded queries). The Non-Occluded Query Masking then masks $Q^{NO}$ to produce masked queries according to their depth $D=[d_{1},d_{2},\cdots,d_{U}]$, leading to the masked queries $Q^{M}=[q_{1}^{M},q_{2}^{M},\cdots,q_{U}^{M}]$. The Query Completion further reconstructs $Q^{M}$ to produce the completed queries $Q^{C}=[q_{1}^{C},q_{2}^{C},\cdots,q_{U}^{C}]$. Finally, the occluded queries $Q^{O}$ and the completed queries $Q^{C}$ are concatenated and fed to the Monocular 3D Detection for 3D detection predictions. Note the inference does not involve the Non-Occluded Query Masking, and it just concatenates the completion of occluded queries $Q^{O}$ (i.e., $Q^{C}$) with the non-occluded queries $Q^{NO}$ and feeds the concatenated queries to the 3D Detection Head for 3D predictions.
@@ 3.3 Non-Occluded Query Masking
Queries predicted by the 3D Backbone are either occluded or non-occluded, depending on whether the corresponding objects are occluded in the input image. In MonoMAE, we mask the non-occluded queries in the feature space to simulate occlusions, aiming to generate pairs of non-occluded and masked (i.e., occluded) queries for learning occlusion-tolerant object representations.
Specifically, we design a Non-Occluded Query Grouping module to identify non-occluded queries and then feed them into a Depth-Aware Masking module to synthesize occlusions, with more detail to be elaborated in the following subsections.
Non-Occluded Query Grouping. The Non-Occluded Query Grouping classifies the queries based on whether their corresponding objects are occluded or non-occluded. With no information about whether the input queries are occluded, we design an occlusion classification network $\mathrm{\Phi}_{\text{O}}$ to predict the occlusion conditions $O^{p}=[o_{1}^{p},o_{2}^{p},\cdots,o_{K}^{p}]$ of queries $Q=[q_{1},q_{2},\cdots,q_{K}]$, where for the $i$-th query $o_{i}^{p}=\mathrm{\Phi}_{\text{O}}(q_{i})$. The Non-Occluded Query Grouping can be formulated by:
@@
where $CE$ is the $Cross$ $Entropy$ loss. We adopted the bipartite matching [4] to match the predicted queries and objects in the image, where only matched queries have ground truth $O^{gt}$ of KITTI 3D [13] about whether they are occluded or not.
Figure 3: Illustration of the Depth-Aware Masking. (a) Objects farther away are usually smaller capturing less visual information. (b) The Depth-Aware Masking determines the mask ratio of an object according to its depth - the closer the object is, the larger the mask ratio is applied, thereby compensating the information deficiency for objects that have larger distances from the camera.
Depth-Aware Masking. We design depth-aware masking to adaptively mask non-occluded query features to simulate occlusions in the feature space, aiming to create non-occluded and occluded (i.e., masked) pairs for learning occlusion-tolerant representations. As illustrated in Figure 3, the depth-aware masking determines the mask ratio according to the object depth - the closer the object, the larger the mask ratio, thereby compensating the information deficiency of distant objects. In addition, we simulate occlusions by masking in the feature space, as masking and reconstructing at the image level is complicated and computationally intensive.
@@
where $q_{i}^{M}$ is the masked query, $q_{i}^{NO}$ is the non-occluded query, and $m_{i}$ is the generated mask.
@@ 3.4 Query Completion
The query completion learns to reconstruct the adaptively masked queries, aiming to produce completed queries whereby the network learns occlusion-tolerant representations that are helpful in detecting occluded objects. We design a completion network $\mathrm{\Phi}_{\text{C}}$ to reconstruct the masked queries. The Completion Network has an hourglass structure consisting of three conv-bn-relu blocks and one conv-bn block for 3D query completion. The completed query $q_{i}^{C}$ is obtained by:
@@
D4LCN [11] CVPR 20 Depth 16.65 11.72 9.51 22.51 16.02 12.55
@@
MonoFlex [68] CVPR 21 None 19.94 13.89 12.07 28.23 19.75 16.89
@@
3.5 Loss Functions
The overall objective consists of three losses including $L_{occ}$, $L_{com}$, and $L_{base}$ where $L_{occ}$ and $L_{com}$ are defined in Equation 2 and Equation 6, and $L_{base}$ denote losses for supervising the 3D box predictions. Specifically, $L_{base}$ includes losses for supervising the 3D box predictions including each object’s 3D locations, height, width, length and orientation. We set the weight for each loss item to 1.0, and the overall loss function is formulated as follows:
$L=L_{occ}+L_{com}+L_{base}.$ (7)
4 Experiments
4.1 Experimental Settings
$\bullet$ NuScenes [3] comprises 1,000 video scenes, including RGB images captured by 6 surround-view cameras. The dataset is split into a training set (700 scenes), a validation set (150 scenes), and a test set (150 scenes). Following [1, 50, 37, 24, 20], the performance on the validation set of nuScenes is reported.
In addition, we perform evaluations on the most representative Car category of KITTI 3D and nuScenes datasets as in prior studies [51, 50, 54, 66]
[…]
Implementation Details. We conduct experiments on one NVIDIA V100 GPU and train the framework for 200 epochs with a batch size of 16 and a learning rate of $2\times 10^{-4}$. We use the AdamW [36] optimizer with weight decay $10^{-4}$. We employ ResNet-50 [16] as the Transformer-based backbone and adopt the 3D detection head from [66] as our detection framework.